# Notions of entropy for unitary representations

### v2

@@ 7.7 Further remarks
Fuglede–Kadison determinants can be hard to compute in general (see [GP14] for some recent examples).
@@
Fuglede–Kadison determinants of matrices over group von Neumann algebras play several roles in connections to other parts of mathematics. A succinct account of some of these is given in [dlH13]. A much more thorough exposition emphasizing $L^{2}$-invariants in topology is given in [Lüc02, Section 3.2]. See also [Lüc02, Chapter 13][KKL25], and the references cited there for an overview of outstanding open problems in this direction, such as Lück’s approximation and determinant conjectures.
In addition, discrete group von Neumann algebras have been generalized to algebras constructed from data such as probability-preserving equivalence relations [FM77a, FM77b] or unimodular random networks [AL07]. See the Introduction and Section 5 of the latter reference for an overview. Lück’s determinant conjecture is generalized accordingly to the ‘measure-theoretic determinant conjecture’ in [LSW10], and shown to have consequences for uniform measure equivalence of discrete groups. Theorem B should generalize in this way as well, for instance by using a version of Example 7.5 over a unimodular random network, but we do not pursue this further here.
@@
Does the construction of a subdiagonal subalgebra from an equivariant random total order shed any light on specific examples of processes on unimodular random networks, or on any of the open questions discussed in [KKL25]?
Hayes’ paper [Hay21a] studies certain measure-preserving systems of algebraic origin for a finitely generated group $\Gamma$ with a left invariant total order. They are constructed from elements of $\mathbf{Z}[\Gamma]$ that he calls ‘lopsided’, which are defined in terms of the order. His main result is that a nondegenerate lopsided group-ring element always gives a factor of a Bernoulli shift.
@@ 8.1 Typical vectors and approximate association
We can classify tuples of vectors in a representation $\pi$ according to their type.
First, for each $k$, let us write $\Sigma_{k}(\mathfrak{A})$ for the set of normalized elements of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. In particular, $\Sigma_{1}(\mathfrak{A})$ is just the state space $\Sigma(\mathfrak{A})$. For each $k$, the set $\Sigma_{k}(\mathfrak{A})$ is a compact and convex subset of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. By the identity (4.5), the pairing isomorphism from (4.4) identifies $\Sigma_{k}(\mathfrak{A})$ with $\Sigma(\mathrm{M}_{k}(\mathfrak{A}))$.
@@
The elements of $\mathcal{X}(\pi,O)$ are the $O$-typical tuples of the representation $\pi$.
We also write
$\Sigma_{k}(\pi):=\big{\{}\Phi^{\pi}_{v_{1},\dots,v_{k}}:\ v_{1},\dots,v_{k}\in H_{\pi}\ \hbox{and}\ \|v_{1}\|^{2}+\cdots+\|v_{k}\|^{2}=k\big{\}}$
and call this the $k$-fold type-set of $\pi$. It is the subset of elements of $\Sigma_{k}(\mathfrak{A})$ that are associated to $\pi$.
@@
for any representation $\pi$ and subset $O$ of $\Sigma_{k}(\mathfrak{A})$.
We often use Definition 8.1 when $O$ is a small neighbourhood around a given ‘target’ completely positive map $\varphi$. In this case we may informally regard elements of $\mathcal{X}(\pi,O)$ as ‘approximately $\varphi$-typical tuples’. This resembles the use of terms such as ‘microstate’ or ‘good model’ in free probability and the study of sofic entropy in ergodic theory: compare [Bow20, Section 2.3], for example.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, let $O$ be any neighbourhood of $\varphi$, and let $\widetilde{O}$ be the corresponding neighbourhood of $\langle\varphi,\cdot\rangle$ under the pairing isomorphism (4.4). Then Lemma 4.4 tells us that
@@
If $\pi$ is a representation and $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, then $\varphi$ is approximately associated to $\pi$ if it lies in the closure of the set of completely positive maps associated to $\pi$. Since $\mathfrak{A}$ is separable, this holds if and only if there is a sequence of $k$-tuples $v_{n,1}$, …, $v_{n,k}$ in $H_{\pi}$ such that
$\Phi^{\pi}_{v_{n,1},\dots,v_{n,k}}(a)\to\varphi(a)\qquad\hbox{for every}\ a\in\mathfrak{A}.$
If $\varphi$ is normalized and approximately associated to $\pi$, and if the sequence of tuples $v_{n,1}$, …, $v_{n,k}$ witnesses the convergence above, then we can normalize those tuples slightly to show that $\pi$ actually lies in the weak closure $\overline{\Sigma_{k}(\pi)}$.
The term ‘approximate association’ is not standard. The literature makes more use of a slightly weaker notion: $\varphi$ is weakly associated to $\pi$ if $\varphi$ can be approximated by finite sums of functionals associated to $\pi$. This is the basis for Godemont’s relation of weak containment for representations: see [Dix77, Section 3.4].
@@
Nevertheless, for some representation the sets of approximately and weakly associated functionals coincide. These include all irreducible representations [Dix77, Theorem 3.4.10], and also regular representations of locally compact groups by a result of Takenouchi (see [Tak55] or [Dix77, Proposition 18.3.5]).
The notion of almost periodic entropy (Definition 8.17 below) involves approximations to a positive functional, but it does not allow for taking finite sums. For this reason, approximate association plays a larger role than weak association in this paper.
@@
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$. The next few lemmas compare typical tuples for these two maps if they are related by association or approximate association. Recall the relation (4.5), and the consequence of this relation that the norm of $\langle\varphi,\cdot\rangle$ is weak continuous when restricted to $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. The next lemma follows at once from this fact and the definition of the relevant topologies.
@@
Fix positive integers $k$ and $\ell$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $a\in\mathrm{M}_{\ell,k}(\mathfrak{A})$, and define a new element of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$ by (4.6): that is,
$\ell\langle\psi,b\rangle:=k\langle\varphi,a^{\ast}ba\rangle\qquad(b\in\mathrm{M}_{\ell}(\mathfrak{A})).$ (8.3)
Then $\psi$ is a continuous function of $(\varphi,a)$, where $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ has the weak topology via the pairing isomorphism (4.4) and $\mathrm{M}_{\ell,k}(\mathfrak{A})$ has the norm topology. ∎
Lemma 8.5.
Let $\psi$ be obtained from $\varphi$ and $a$ as in (8.3). For any neighbourhood $U$ of $\psi$, there is a neighbourhood $V$ of $\varphi$ such that
$\pi^{(\ell,k)}(a)[\mathcal{X}(\pi,V)]\subset\mathcal{X}(\pi,U)$
@@
Consider a $k$-tuple $v_{1}$, …, $v_{k}$ in $H_{\pi}$, and define the $\ell$-tuple $y_{1}$, …, $y_{\ell}$ by
$[y_{1},\dots,y_{\ell}]^{\mathrm{T}}:=\pi^{(\ell,k)}(a)[v_{1},\dots,v_{k}]^{\mathrm{T}}.$
If these two tuples have respective types $\varphi^{\prime}$ and $\psi^{\prime}$, then Lemma 4.5 tells us that $\psi^{\prime}$ is again related to $\varphi^{\prime}$ as in (8.3). Now the existence of $V$ for a given $U$ follows by the continuity given by Lemma 8.4. ∎
Corollary 8.6.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$, and assume that $\psi$ is approximately associated to $\pi_{\varphi}$. Then for any neighbourhood $U$ of $\psi$ there is a neighbourhood $V$ of $\varphi$ such that
$\mathcal{X}(\pi,V)\neq\emptyset\quad\Rightarrow\quad\mathcal{X}(\pi,U)\neq\emptyset.$
for any representation $\pi$.
Proof.
Let $\varphi$ be associated to $\pi_{\varphi}$ by the cyclic tuple $x_{1}$, …, $x_{k}$. By cyclicity and Lemma 4.2, there is some $a\in\mathrm{M}_{\ell,k}(\mathfrak{A})$ such that the tuple defined by
$[y_{1},\dots,y_{\ell}]^{\mathrm{T}}:=\pi^{(\ell,k)}(a)[x_{1},\dots,x_{k}]^{\mathrm{T}}$
satisfies $\psi^{\prime}:=\Phi^{\pi}_{y_{1},\dots,y_{\ell}}\in U$. Now apply Lemma 8.5 to $\varphi$, $\psi^{\prime}$ and $a$. ∎
Lemma 8.5 has the following variant, which we formulate separately for later reference. We prove it separately as well, but it is actually a special case of Lemma 8.5 by taking $a=1\otimes Q$ for some complex matrix $Q$.
Lemma 8.7.
Let $\varphi\in\mathfrak{B}(\mathfrak{A};\mathrm{M}_{k})_{+}$, let $Q\in\mathrm{M}_{k,\ell}$, and define
$\psi(b):=Q^{\ast}\varphi(b)Q\qquad(b\in\mathfrak{A}).$ (8.4)
For any neighbourhood $U$ of $\psi$, there is a neighbourhood $V$ of $\varphi$ such that
$\big{\{}([x_{1},\dots,x_{k}]Q)^{\mathrm{T}}:\ [x_{1},\dots,x_{k}]^{\mathrm{T}}\in\mathcal{X}(\pi,V)\big{\}}\subset\mathcal{X}(\pi,U)$
for any representation $\pi$.
Proof.
First consider any representation $\pi$ and $k$-tuple $V=[x_{1},\dots,x_{k}]$ in $H_{\pi}$. Define an $\ell$-tuple in the same space by $[y_{1},\dots,y_{\ell}]:=VQ$. Then
$\Phi^{\pi}_{y_{1},\dots,y_{k}}(b)=Q^{\ast}V^{\ast}\pi(b)VQ=Q^{\ast}\cdot\Phi^{\pi}_{x_{1},\dots,x_{k}}(b)\cdot Q\qquad(b\in\mathfrak{A}).$
This is the same relationship between completely positive maps as in (8.4). The result follows because, with $Q$ fixed, that relationship gives $\psi$ as a continuous function of $\varphi$. ∎
8.2 Pairs and sums of typical vectors
@@
We borrow the next term from Furstenberg’s classic work [Fur67] in ergodic theory.
Definition 8.8.
In particular, the diagonal joining is the joining defined by
Notice that, if $\varphi\in\Sigma_{k}(\mathfrak{A})$ and $\psi\in\Sigma_{\ell}(\mathfrak{A})$, then any joining of them lies in $\Sigma_{k+\ell}(\mathfrak{A})$.
If $\varphi$ and $\psi$ are disjoint, then $\pi_{\varphi}$ and $\pi_{\psi}$ must appear orthogonally in any representation that contains both of them. Therefore $\mathrm{diag}(\varphi,\psi)$ is the only joining of $\varphi$ and $\psi$. The reverse implication holds as well, because any common subrepresentation can be used to construct a non-diagonal joining. (In fact, starting with Furstenberg’s paper [Fur67], the uniqueness of the diagonal joining is taken as the definition of ‘disjointness’ in ergodic theory, where the lack of orthogonal complements makes other definitions impractical or senseless.)
The next lemma is a robust form of this statement.
@@
If $\varphi$ and $\psi$ are disjoint and $O$ is a neighbourhood of $\mathrm{diag}(\varphi,\psi)$, then there are neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$ such that the following holds:
@@
The result is trivial if either $\varphi$ and $\psi$ is zero, so assume thay are both nonzero. Let $a:=\overline{\mathrm{tr}}_{k}\varphi(1)$ and $b:=\overline{\mathrm{tr}}_{\ell}\psi(1)$, so these are both positive, and consider normalizing $\varphi$ to $\varphi/a$ and $\psi$ to $\psi/b$. On elements of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k+\ell})_{+}$, the corresponding normalization is
$\theta\mapsto\mathrm{diag}(\sqrt{a}I_{k},\sqrt{b}I_{\ell})^{-1/2}\cdot\theta\cdot\mathrm{diag}(\sqrt{a}I_{k},\sqrt{b}I_{\ell})^{-1/2}.$
This normalization reduces our work to the case when $\varphi\in\Sigma_{k}(\mathfrak{A})$ and $\psi\in\Sigma_{\ell}(\mathfrak{A})$.
Since $\mathfrak{A}$ is separable, the relevant topologies are first countable. Towards a contradiction, we may therefore suppose that there is a sequence $(\theta^{\prime}_{n})_{n\geq 1}$ in $\Sigma_{k+\ell}(\mathfrak{A})$ satisfying
$(\theta^{\prime}_{n})[K]\to\varphi\quad\hbox{and}\quad(\theta^{\prime}_{n})[L]\to\psi\qquad\hbox{but}\qquad\theta^{\prime}_{n}\not\to\theta.$
Since $\Sigma_{k+\ell}(\mathfrak{A})$ is compact, we may pass to a subsequence so that $\theta^{\prime}_{n}$ actually converges to a different element $\theta^{\prime}$ of $\mathfrak{B}(\mathfrak{A};\mathrm{M}_{k+\ell})_{+}$. This is now a non-diagonal joining of $\varphi$ and $\psi$, contradicting their disjointness.
@@
The assumption of disjointness in Lemma 8.9 is necessary. Indeed, if $\pi_{\varphi}$ is finite-dimensional and irreducible, then $\mathcal{X}(\pi,U)$ is nonempty for any neighourhood $U$ of $\varphi$, but $\mathrm{diag}(\varphi,\varphi)$ is not necessarily approximately associated to $\pi$, only to $\pi^{\oplus 2}$, by similar reasoning to Example (LABEL:eq:sigma-not-convex).
The last two results in this subsection concern typical vectors for the sum $\gamma:=\varphi+\psi$. Here we restrict our attention to the case $k=\ell=1$. The case $k=\ell>1$ would not involve any new ideas, but would require heavier notation. Later in the paper we use the constructions of Subsection 4.3 to avoid needing that case.
@@
$(\pi(a)x,\pi(1-a)x)\in\mathcal{X}(\pi,U)\times\mathcal{X}(\pi,V)\qquad\hbox{for all}\qquad x\in\mathcal{X}(\pi,W)$
@@
The representations $\pi_{\varphi}$ and $\pi_{\psi}$ are disjoint by assumption, and they are both contained in $\pi_{\gamma}$ by Proposition 3.2. Therefore we may identify $\pi_{\gamma}$ with ${\pi_{\varphi}\oplus\pi_{\psi}}$, and the resulting orthogonal projection $P$ from $H_{\gamma}$ to $H_{\varphi}$ lies in the centre of $\pi_{\gamma}(\mathfrak{A})^{\prime\prime}$ by Lemma 3.1.
Now suppose that $\gamma$ is associated to $\pi_{\gamma}$ by the cyclic vector $v$. The Kaplansky density theorem [Dix81, Section I.3.5] applied to $\pi_{\gamma}(\mathfrak{A})$ gives an element $a\in\mathfrak{A}$ such that $0\leq a\leq 1$ and such that $\pi_{\gamma}(a)v$ lies as close as we wish to $Pv$. In particular, we may choose $a$ so that
@@
First let $W$ be a neighbourhood of $\gamma$. Given $x$, $y\in H_{\pi}$, observe that
@@
8.3 The strong quotient topology on representations
Approximate association leads to a natural equivalence relation on representations, and an associated topology on equivalence classes.
First, if $X$ is any topological space, then its hypersace $\mathcal{K}(X)$ is the set of its nonempty compact subsets.
Given a separable representation $\pi$, let us call $\overline{\Sigma_{k}(\pi)}$ the $k$-summary of $\pi$, so this is an element of $\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$. We refer to the whole sequence
$\overline{\Sigma_{\bullet}(\pi)}:=(\overline{\Sigma_{1}(\pi)},\overline{\Sigma_{2}(\pi)},\dots)\in\prod_{k=1}^{\infty}\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$ (8.5)

### v3

@@ 7.7 Further remarks
Fuglede–Kadison determinants can be hard to compute in general (see [70] for some recent examples).
@@
Fuglede–Kadison determinants of matrices over group von Neumann algebras appear in several connections with other parts of mathematics. A succinct account of some of these is given in [42]. A much more thorough exposition emphasizing $L^{2}$-invariants in topology is given in [107, Section 3.2]. See also [107, Chapter 13][97], and the references cited there for an overview of outstanding open problems in this direction, such as Lück’s approximation and determinant conjectures.
In addition, discrete group von Neumann algebras have been generalized to algebras constructed from data such as probability-preserving equivalence relations [56, 57] or unimodular random networks [3]. See the Introduction and Section 5 of the latter reference for an overview. Lück’s determinant conjecture is generalized accordingly to the ‘measure-theoretic determinant conjecture’ in [108], and shown to have consequences for uniform measure equivalence of discrete groups. I expect that Theorem B generalizes in this way as well, for instance by using a version of Example 7.5 over a unimodular random network.
@@
Does the construction of a subdiagonal subalgebra from an equivariant random total order shed any light on specific examples of processes on unimodular random networks, or on any of the open questions discussed in [97]?
Hayes’ paper [83] studies certain measure-preserving systems of algebraic origin for a finitely generated group $\Gamma$ with a left invariant total order. They are constructed from elements of $\mathbf{Z}[\Gamma]$ that he calls ‘lopsided’, which are defined in terms of the order. His main result is that a nondegenerate lopsided group-ring element always gives a factor of a Bernoulli shift.
@@ 8.1 Typical vectors and approximate association
Consider again a general separable, unital C*-algebra $\mathfrak{A}$. We can classify tuples of vectors in a representation $\pi$ of $\mathfrak{A}$ according to their type.
For each $k$, let $\Sigma_{k}(\mathfrak{A})$ be the set of normalized elements of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. In particular, $\Sigma_{1}(\mathfrak{A})$ is just the state space $\Sigma(\mathfrak{A})$. By the identity (4.9), the pairing isomorphism from (4.8) identifies $\Sigma_{k}(\mathfrak{A})$ with $\Sigma(\mathrm{M}_{k}(\mathfrak{A}))$, and this is a norm-bounded subset of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})$. It is also compact and convex by the Banach–Alaoglu theorem.
@@
The elements of $\mathcal{X}(\pi,O)$ are the $O$-typical tuples of the representation $\pi$. In addition, let
$\Sigma_{k}(\pi):=\big{\{}\Phi^{\pi}_{v_{1},\dots,v_{k}}:\ v_{1},\dots,v_{k}\in H_{\pi}\ \hbox{and}\ \|v_{1}\|^{2}+\cdots+\|v_{k}\|^{2}=k\big{\}}.$
This is the subset of elements of $\Sigma_{k}(\mathfrak{A})$ that are associated to $\pi$.
@@
for any representation $\pi$ and any $O\subset\Sigma_{k}(\mathfrak{A})$.
We often use Definition 8.1 when $O$ is a small neighbourhood around a given ‘target’ completely positive map $\varphi$. In this case we may informally describe elements of $\mathcal{X}(\pi,O)$ as ‘approximately $\varphi$-typical tuples’. This resembles the use of terms such as ‘microstate’ in free probability or ‘good model’ in the study of sofic entropy in ergodic theory: compare [26, Section 2.3], for example.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, let $O$ be a neighbourhood of $\varphi$, and let $\widetilde{O}$ be the corresponding neighbourhood of $\langle\varphi,\cdot\rangle$ under the pairing isomorphism (4.8). Then Lemma 4.7 tells us that
@@
If $\pi$ is a representation and $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, then $\varphi$ is approximately associated to $\pi$ if it lies in the closure of the set of completely positive maps associated to $\pi$. Because of the second countability from Lemma 4.8, this holds if and only if there is a sequence of $k$-tuples $v_{n,1}$, …, $v_{n,k}$ in $H_{\pi}$ such that
$\Phi^{\pi}_{v_{n,1},\dots,v_{n,k}}(a)\to\varphi(a)\qquad\hbox{for every}\ a\in\mathfrak{A}\ \hbox{as}\ n\to\infty.$ (8.3)
If $\varphi$ is approximately associated to $\pi$ and also normalized, and if the tuples $v_{n,1}$, …, $v_{n,k}$ witness the convergence in (8.3), then we can normalize those tuples slightly to show that $\pi$ actually lies in the weak closure $\overline{\Sigma_{k}(\pi)}$.
The term ‘approximate association’ is not standard. The literature makes more use of a weaker notion: $\varphi$ is weakly associated to $\pi$ if $\varphi$ can be approximated by finite sums of maps associated to $\pi$. This is the basis for Godemont and Fell’s relation of weak containment for representations: see [73][59] or [48, Section 3.4].
@@
Nevertheless, for some representations the sets of approximately and weakly associated functionals coincide. These include all irreducible representations [48, Theorem 3.4.10], and also regular representations of locally compact groups by a result of Takenouchi (see [140] or [48, Proposition 18.3.5]). For a discrete group, the latter result can actually be deduced from Proposition 3.5. In that case, if $\lambda$ is the regular representation, then it has a tracial cyclic vector, and so Proposition 3.5 shows that $\Sigma_{1}(\lambda^{\oplus\infty})$ is actually equal to $\Sigma_{1}(\lambda)$. Therefore all the intermediate spaces $\Sigma_{1}(\lambda^{\oplus k})$ are equal to $\Sigma_{1}(\lambda)$ as well. On the other hand, given an element of $\Sigma_{1}(\lambda^{\oplus m})$ and an element of $\Sigma_{1}(\lambda^{\oplus n})$, any convex combination of them lies in $\Sigma_{1}(\lambda^{\oplus(m+n)})$.
Our notion of almost periodic entropy (Definition 9.4 below) involves approximations to a positive functional, but it does not allow for taking finite sums. For this reason, approximate association plays a larger role than weak association in the present paper.
@@
In the next lemma and corollary, we compare typical tuples for two completely positive maps if they are related by association or approximate association. First, fix an $\ell$-by-$k$ matrix $[a_{ij}]$ of elements of $\mathfrak{A}$. If $\pi$ is a representation and ${v_{1},\dots,v_{k}}\in H_{\pi}$, recall that we can define a new $\ell$-tuple ${y_{1},\dots,y_{\ell}}$ in $H_{\pi}$ using $[a_{ij}]$ and ${v_{1},\dots,v_{k}}$ as in formula (4.2). The resulting type of ${y_{1},\dots,y_{\ell}}$ is then related to the type of ${v_{1},\dots,v_{k}}$ by Lemma 4.3. Because of the weak continuity given by that lemma, we obtain the following.
@@
Let $\psi$ be obtained from $\varphi$ and $[a_{ij}]$ as in Lemma 4.3. For any neighbourhood $O$ of $\psi$, there is a neighbourhood $U$ of $\varphi$ such that
$\big{\{}[\pi(a_{ij})]\cdot[v_{1},\dots,v_{k}]^{\mathrm{T}}:\ [v_{1},\dots,v_{k}]^{\mathrm{T}}\in\mathcal{X}(\pi,U)\big{\}}\subset\mathcal{X}(\pi,O)$
for any representation $\pi$.
In particular, suppose that
$\psi(b)=(Q^{\mathrm{T}})^{\ast}\varphi(b)Q^{\mathrm{T}}\qquad(b\in\mathfrak{A})$
for some $Q\in\mathrm{M}_{\ell,k}$, as in (4.5). Then, for any neighbourhood $O$ of $\psi$, there is a neighbourhood $U$ of $\varphi$ such that
$(I_{H_{\pi}}\otimes Q)[\mathcal{X}(\pi,U)]\subset\mathcal{X}(\pi,O)$
for any representation $\pi$ (identifying $H_{\pi}^{\oplus k}$ with $H_{\pi}\otimes\mathbf{C}^{\oplus k}$ as in (4.5)). ∎
Corollary 8.5.
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$, and assume that $\psi$ is approximately associated to $\pi_{\varphi}$. Then for any neighbourhood $O$ of $\psi$ there is a neighbourhood $U$ of $\varphi$ such that
$\mathcal{X}(\pi,U)\neq\emptyset\quad\Rightarrow\quad\mathcal{X}(\pi,O)\neq\emptyset.$
@@
Let $\varphi$ be associated to $\pi_{\varphi}$ by the cyclic tuple $x_{1}$, …, $x_{k}$. By cyclicity and Lemma 4.2, there is some $\ell$-by-$k$ matrix $[a_{ij}]$ of elements of $\mathfrak{A}$ such that the tuple defined by
$[y_{1},\dots,y_{\ell}]^{\mathrm{T}}:=[\pi(a_{ij})]\cdot[x_{1},\dots,x_{k}]^{\mathrm{T}}$
satisfies $\psi^{\prime}:=\Phi^{\pi}_{y_{1},\dots,y_{\ell}}\in O$. Now apply Lemma 8.4 to $\varphi$, $\psi^{\prime}$ and $[a_{ij}]$. ∎
8.2 Pairs and sums of typical tuples
@@
Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ and $\psi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{\ell})_{+}$. The next lemma is a robust form of Lemma 4.5.
@@
If $\varphi$ and $\psi$ are disjoint, then for every neighbourhood $O$ of $\mathrm{diag}(\varphi,\psi)$ there are neighbourhoods $U$ of $\varphi$ and $V$ of $\psi$ such that the following holds:
@@
We prove the first conclusion by contraposition. Assume that $\mathrm{diag}(\varphi,\psi)$ has a neighbourhood $O$ for which no pair of neighbourhoods $U$ and $V$ gives the desired implication. By the second countability from Lemma 4.8, we may therefore choose a sequence $(\theta_{n})_{n\geq 1}$ in $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}\setminus O$ such that
$\theta_{n}[K]\to\varphi\qquad\hbox{and}\qquad\theta_{n}[L]\to\psi.$
This sequence $(\theta_{n})_{n\geq 1}$ must be uniformly bounded in the dual norm because of the identity (4.9) and the relation
$\overline{\mathrm{tr}}_{k+\ell}\theta_{n}(1)=\frac{k}{k+\ell}\overline{\mathrm{tr}}_{k}\big{(}\theta_{n}(1)[K]\big{)}+\frac{\ell}{k+\ell}\overline{\mathrm{tr}}_{\ell}\big{(}\theta_{n}(1)[L]\big{)}.$
Therefore, by the Banach–Alaoglu theorem, $(\theta_{n})_{n\geq 1}$ has a subsequential limit in the weak topology. This limit must be a joining of $\varphi$ and $\psi$, but also it cannot lie in $O$ and so it must be different from $\mathrm{diag}(\varphi,\psi)$. Therefore $\varphi$ and $\psi$ are not disjoint.
@@
The assumption of disjointness in Lemma 8.6 is necessary. Indeed, if $\pi_{\varphi}$ is finite-dimensional and irreducible, then $\mathcal{X}(\pi,U)$ is nonempty for any neighourhood $U$ of $\varphi$, but $\mathrm{diag}(\varphi,\varphi)$ is not necessarily approximately associated to $\pi$, only to $\pi^{\oplus 2}$, by reasoning similar to Example 8.2.
Now we consider typical vectors for the sum $\gamma:=\varphi+\psi$. Here we restrict our attention to the case $k=\ell=1$. The case $k=\ell>1$ would not involve any new ideas, but would require heavier notation. Later in the paper we use the constructions of Subsection 4.3 to avoid needing that case.
@@
$\{(\pi(a)x,\pi(1-a)x):\ x\in\mathcal{X}(\pi,W)\}\subset\mathcal{X}(\pi,U)\times\mathcal{X}(\pi,V)$
@@
The representations $\pi_{\varphi}$ and $\pi_{\psi}$ are disjoint by assumption, and they are both contained in $\pi_{\gamma}$ by Proposition 3.2. Therefore, by the uniqueness of GNS representations, we may identify $\pi_{\gamma}$ with ${\pi_{\varphi}\oplus\pi_{\psi}}$. The resulting orthogonal projection $P$ from $H_{\gamma}$ to $H_{\varphi}$ lies in the centre of $\pi_{\gamma}(\mathfrak{A})^{\prime\prime}$ by Lemma 3.1.
Now suppose that $\gamma$ is associated to $\pi_{\gamma}$ by the cyclic vector $v$. The Kaplansky density theorem [49, Section I.3.5] applied to $\pi_{\gamma}(\mathfrak{A})$ gives an element $a\in\mathfrak{A}$ such that $0\leq a\leq 1$ and such that $\pi_{\gamma}(a)v$ lies as close as we wish to $Pv$. In particular, we may choose $a$ so that
@@
First let $W$ be a neighbourhood of $\gamma$. Given a representation $\pi$ and $x$, $y\in H_{\pi}$, observe that
@@
8.3 The strong-quotient topology on representations
Approximate association leads to a natural equivalence relation on representations, and to a topology on equivalence classes.
First, if $X$ is any topological space, then its hyperspace $\mathcal{K}(X)$ is the set of its nonempty compact subsets.
Given now a separable representation $\pi$, let us call $\overline{\Sigma_{k}(\pi)}$ the $k$-summary of $\pi$. It is an element of $\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$. We refer to the whole sequence
$\overline{\Sigma_{\bullet}(\pi)}:=\big{(}\overline{\Sigma_{1}(\pi)},\overline{\Sigma_{2}(\pi)},\dots\big{)}\in\prod_{k=1}^{\infty}\mathcal{K}(\Sigma_{k}(\mathfrak{A}))$ (8.4)